OpenStudy (anonymous):

The radius of a cone-shaped tank is 4 feet less than its height. If the height of the tank is (x - 3) feet, the expression below shows the volume of the tank. 1 over 3π(x - 7)2 (x - 3) What does the factor π (x - 7)2 represent? The total surface area of the tank The area of the curved sides of the tank The area of the circular base of the tank The area of the base of about three such tanks

OpenStudy (mathstudent55):

Do you know the formula for the volume of a cone?

OpenStudy (anonymous):

well i forgot

OpenStudy (mathstudent55):

Since the problem mentions a cone and the volume of a cone, it will be very helpful to refresh your memory.

OpenStudy (mathstudent55):

The volume of a cone is \(V = \dfrac{1}{3} \pi r^2 h\) where V = volume of cone, r = radius of the base h = height of the cone

OpenStudy (mathstudent55):

Now let's look at your cone. For your cone, we are told h = x - 3 We are also told the radius is 4 less than the height. What is 4 less than x - 3?

OpenStudy (anonymous):

so is the answer c.

OpenStudy (mathstudent55):

Exactly. Since the height is x - 3, the radius is x - 7. The area of a circle is \(A = \pi r^2\) Here you have \(A = \pi (x - 7)^2\) That is the area of the base.

OpenStudy (anonymous):

yayyy i got it right and thanks for the medal :) and for helping me here you go.

OpenStudy (mathstudent55):

You're welcome.